OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)?

OpenStudy (anonymous):

\[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\] use tis eq in your case x1=-2 y1=0 ,x2=2, y2=8 simply put values in eq.you will get the resuly..